The perimeter of a rectangle is 250 yards. The length is ten less than the width. The measurements are:
Answer by algebrahouse.com(1659)

x = width
x - 10 = length {length is 10 less than the width}

Perimeter of a rectangle is 2(width) + 2(length)

2x + 2(x - 10) = 250 {perimeter is 2(width) + 2(length)}
2x + 2x - 20 = 250 {used distributive property}
4x - 20 = 250 {combined like terms}
4x = 270 {added 20 to each side}
x = 67.5 {divided each side by 4}
x - 10 = 57.5 {substituted 67.5, in for x, into x - 10}

width = 67.5 yds
length = 57.5 yds

For more help from me, visit: www.algebrahouse.com